# MRNaB: Mixed Reality-based Robot Navigation Interface using Optical-see-through MR-beacon

### v1

@@ Abstract
Recent advancements in robotics have led to the development of numerous interfaces to enhance the intuitiveness of robot navigation. However, the reliance on traditional 2D displays imposes limitations on the simultaneous visualization of information. Mixed Reality (MR) technology addresses this issue by enhancing the dimensionality of information visualization, allowing users to perceive multiple pieces of information concurrently. This paper proposes Mixed reality-based robot navigation interface using an optical-see-through MR-beacon (MRNaB), a novel approach that incorporates an MR-beacon, situated atop the real-world environment, to function as a signal transmitter for robot navigation. This MR-beacon is designed to be persistent, eliminating the need for repeated navigation inputs for the same location. Our system is mainly constructed into four primary functions: ”Add”, ”Move”, ”Delete”, and ”Select”. These allow for the addition of a MR-beacon, location movement, its deletion, and the selection of MR-beacon for navigation purposes, respectively. The effectiveness of the proposed method was then validated through experiments by comparing it with the traditional 2D system. As the result, MRNaB was proven to increase the performance of the user when doing navigation to a certain place subjectively and objectively. For additional material, please check: https://mertcookimg.github.io/mrnab
@@ I Introduction
Recently, many researchers have focused on mixed reality (MR) [7, 8, 9] as a tool for enhancing human-computer interaction by offering a more intuitive and immersive interface. MR significantly improves the user experience by providing a unified view of virtual and real-world elements, thereby minimizing the need for attention shifts and enhancing operational effectiveness[10, 11, 12]. Numerous studies have introduced innovative robot navigation interfaces through the use of an MR device called Hololens 2 [6, 13, 14, 15]. However, specifying a navigation destination within a specific location, especially during delivery processes in households or companies, can be challenging. Existing interfaces encounter limitations in two main areas: Firstly, existing interfaces set the robot’s destination with markers such as an arrow. With only an arrow, users often find it difficult to visualize the robot status on the destination, especially in the space-constrained area. This difficulty can cause the user to fail to navigate the robot to its intended location because users can’t visualize the robot’s pose at the destination. Secondly, users need to repeatedly specify the robot’s pose for frequently visited destinations, which diminishes efficiency.
This paper introduces MRNaB, a novel approach for robotic navigation by employing a persistent MR-beacon that mirrors the actual shape of the robot and retains its position even after the project is restarted. Fig. 1 shows the concept of the system. This approach not only simplifies the task of visualizing the robot’s destination in navigation but also streamlines the process of setting destinations that are regularly used. In order to implement the system, we provide the user with 4 main functions which are ”add” to add the MR-beacon from the floor, ”move” to move the location of the MR-beacon, ”select” to select the MR-beacon for the robot navigation, and ”delete” to delete the MR-beacon that has been made. We also implemented a database system, which eliminates the need for users to repeatedly specify the robot’s location. This feature is particularly beneficial for environments that require repetitive navigation to the same place, such as households or companies, especially for delivery tasks. Our contributions are summarized as follows:
  • This paper proposed MRNaB, a mixed reality navigation interface with MR-beacon shaped like the robot, improving destination visualization and navigation efficiency.
@@
This paper consists of six sections including this one. Section II explains the Related works. Sections III and IV introduce Proposed System Design and System Implementation. In Section V, experimental results are shown to confirm the usefulness of the proposed method. Section VI concludes this paper.
@@ II Related Works
$\bullet$ Real World Spatial Understanding: A 2D map, generated by SLAM using the robot’s Lidar, provides only a rough estimation of the environment based on the Lidar’s height. This limitation becomes apparent in typical household structures, which do not conform to uniform shapes like cubes, thus complicating spatial comprehension from a 2D perspective. 3D maps, though can give the understanding of the space, but is not as clear as the real-world itself and the creation of 3D maps is resource-intensive, demanding significant computational power and memory, especially for dynamic environments with frequent changes.
Our research aims to bridge the identified challenges by augmenting user visual capabilities through the integration of real-world information and 2D maps generated by SLAM for robotic navigation. This approach addresses several key issues: it enables immediate impact on real-world robots, mitigates resource constraints, adapts effectively to dynamic environments, and enhances spatial understanding, and eliminates the need for users to constantly shift their focus between the display and the real-world environment.
@@ II-B MR Robot Navigation Interface
In the field of MR for robotic navigation, research initiatives have explored innovative methods to enhance human-robot interaction. Studies include using Hololens 2 to visualize robot information through occlusions [24] and employing markers and signs to indicate robot motion intent, safety distances, and interaction guidelines [25, 26]. Specialized interfaces for wheelchair robots have been developed to display intent [27, 28], and multi-user interfaces facilitate simultaneous interactions with multiple robots [29]. Lee et al. also explore the eye gaze field by using it to operate the robot manually [13]. Kot et al. combine a joystick and Hololens 2 to operate the robot manually from other places [30]. However, these studies are for manual navigation or don’t directly navigate the robot.
This paper focuses on the MR interface for autonomous navigation. For autonomous navigation, one notable example is ARviz [6], which enhances 2D maps created by SLAM through the use of Hololens 2, enabling visualization of information provided by Rviz. Robot navigation is facilitated by arrows created by ”air tap” hand gestures to create the goal of the navigation. Some also use waypoints for navigation goals [31, 32, 33]. Chen et al. made an approach involving an interface that employs holograms and ”drag and drop” interface for robot information visualization and navigation by utilizing 3D map created by SLAM [14]. Wu et al. use the 2D map with waypoints interface to navigate the robot [34]. Zhang et al. also explore the utilization of eye gaze and head movement for robot navigation and robot arm control [15].
@@
Destination Status Visualization
@@
$\bullet$ Destination Status Visualization: Users are unable to visualize the actual condition or status of the robot’s destination. This limitation makes navigation difficult, especially in space-constrained areas.
@@
Our research proposes to overcome these challenges by enhancing the navigation interface by representing the robot’s goal pose with an object that matches the robot’s actual dimensions, thus making it easier for user to visualize where their robot wants to go. Our research also proposes implementing a database system that allows users to retrieve previously set destinations upon re-running the program. To sum up, the comparison between our interface and other interfaces can be expressed in Table I. 2D interface despite having destination status visualization and destination persistency, has a problem with seamless attention and real-world spatial understanding since it only uses the map. MR interface addresses this problem. However not having the ability to visualize the robot’s destination and a feature to navigate to repetitive places makes navigation difficult. This research tries to bridge all these challenges by providing all the features provided by both 2D interfaces and other MR interfaces.
Figure 2: System Design Diagram. The user interacts with Hololens 2 by using gestures to access the hand menu and interact with the MR beacon to access the beacon functions. Our system also has a database that saves the information (Pose and ID) of the robot and uses it during the first co-localization. Hololens 2 receives robot data while sending the log and goal pose to the ROS 2 side. ROS 2 has a log recorder and TF converter to convert robot and map information to only robot information. Goal pose will be sent directly using the Kachaka gprc ROS 2 bridge to the ”Kachaka” robot.
@@ III System Design
In order to access the MR-beacon to navigate the robot, the hand menu is used by doing the ”Hand Constraint Palm Up” movement. Hand menu is shown in Fig. 3. In the hand menu, there will be beacon button and stage button. Stage button is only used for experiment. Beacon button will then further be divided into 6 buttons which are the back button, off button, add button, move button, select button, and delete button. Back button is used to return to the main menu and off button is used to remove all the functionality of the MR-beacon. The main functions for this system are add button, move button, select button, and delete button.
III-A Add Button
Add button is used to generate MR-beacon from the floor. After pressing this button, user will enter ”Add Mode”. Fig. 4 shows the ”Add Mode” process for MR-beacon.
Initially, there is no MR-beacon on the floor. When user uses ”Air Tap” gesture (pointer down) to the floor, beacon will be generated from the floor and follow the location of the pointer of the user by dragging the pointer. In order to fixate the location of the MR-beacon, user needs to release the ”Air Tap” and after that direction setting will start where MR-beacon direction will follow the location of the pointer of the user. To fixate the direction of the MR-beacon, user needs to do ”Air Tap” gesture once more (pointer click).
III-B Move Button
Move button is used to move or adjust the pose of the MR-beacon to other poses. After pressing this button, user will enter ”Move Mode”, and user can move the MR-beacon that was generated before. Fig. 5 shows ”Move Mode” process which works similarly to ”Add Mode”. However, instead of doing the ”air tap” movement to the floor, the user has to do ”air tap” to the MR-beacon.
The initial state starts with MR-beacon already on the floor. When user does ”Air Tap” gesture to the MR-beacon, the MR-beacon will turn yellow and user can drag the location of the MR-beacon freely on the floor. In order to fixate the location of the MR-beacon, user needs to release the ”Air Tap” and after that direction setting will start where MR-beacon direction will follow the location of the pointer of the user. In order to fixate the direction of the MR-beacon, user needs to do ”Air Tap” gesture once more (pointer click).
@@ III-B Move Button
Figure 4: Add Mode Process. (a) shows the initial state where there is no MR-beacon on the floor. (b) When user does the ”air tap” movement to the floor, MR-beacon will be generated and location setting starts. (c) As user drags the pointer, MR-beacon will follow the pointer. Once ”air tap” is released, MR-beacon will fix the location and start the direction setting. (d) To fix the direction, user has to do one more ”air tap”.
@@
Figure 5: Move Mode Process. (a) shows the initial state where there an MR-beacon on the floor. (b) When user does the ”air tap” movement to the MR-beacon, MR-beacon will be generated and the location setting starts. (c) As user drags the pointer, MR-beacon will follow the pointer. Once ”air tap” is released, MR-beacon will fix the location and start the direction setting. (d) To fix the direction, user has to do one more ”air tap”.
@@
Figure 6: Select Mode Process. (a) shows the initial state where there is already an MR-beacon on the floor. (b) By clicking the MR-beacon, MR-beacon’s location will be sent to the real robot. (c) Navigation will be started once robot receives the information. (d) Robot reaches the location of the MR-beacon.
@@
III-C Select Button
Select button is used to send the pose of the MR-beacon to the real robot. After pressing this button, user will enter ”Select Mode”, and user can navigate the robot to any of the MR-beacon generated. ”Select Mode” process is shown in Fig. 6. Initial state starts with MR-beacon on the floor already. By doing ”Air Tap” gesture (pointer click) to the MR-beacon, the pose of the MR-beacon will be transformed to be relative to the QR code and be sent to the robot. Robot will then do navigation to the targeted pose.
III-D Delete Button
Delete button is used to delete MR-beacon. After pressing this button, user will enter ”Delete Mode” where user can delete any generated MR-beacon on the floor. ”Delete Mode” process is shown in Fig. 7. Initial state starts with MR-beacon on the floor. By doing ”Air Tap” gesture (pointer click) to any MR-beacon, MR-beacon will be deleted from the floor. The information of that MR-beacon will also be deleted from the database.
@@ IV-B Co-localization
In order to align the coordinate of ROS and Hololens 2 (Unity), co-localization is required. In this system, we use the same co-localization process used by ARviz [6] by using vuforia. By using vuforia, real object’s prototype will first be put to the vuforia database and then the object will be accessed as a unity object. All of the object made in the unity system will have the parent set to this QR code. We put this QR code to be on the same location of the map location in ROS system, in order to make the transformation from Unity coordinate system to ROS coordinate system.
IV-C Database
@@
The database is used to store the information about the MR-beacon. All the information of the MR-beacon that will be stored includes: First, the name of the robot which also can be used as the ID. Second position of the MR-beacon is represented by $x$, $y$, $z$ for $x$-axis, $y$-axis, and $z$-axis of the location of the robot relative to the QR code respectively. Third, rotation of the MR-beacon represented by $q_{x}$, $q_{y}$, $q_{z}$, and $q_{w}$ quaternion of the robot relative to the QR code respectively. For ID, we used Globally Unique Identifier (GUID) to make it unique. After the direction setting in the add function, the database will store the information of the established MR-beacon as new data. After the direction setting in the move function, the database will check the moved MR-beacon ID to update the information of the pose. After the delete function is used, the database will remove the item from the database based on the ID. Lastly when the user starts the project, if the QR code has been found, all of the stored beacons will be instantiated. The database’s pseudocode is shown in Algorithm 1.
IV-D MR-beacon
IV-D1 MR-beacon Design
The dimensions of the MR-beacon are designed to mirror those of the real robot we used, as illustrated in Fig. 8. While the actual shape of the robot is not precisely a block, we simplified its form for the MR-beacon’s design, disregarding specific elements like the wheels and the front camera.
IV-D2 Add Function
For MR-beacon add function, since there is no MR-beacon in the first place, we put a transparent plane object at the top of the floor whose origin is located exactly at the QR Code location. To this object, we then put object manipulator provided by MRTK, while restricting the movement to only $x$-axis and $z$-axis, as well as restricting the rotation in all directions. Once user starts doing the ”air tap” movement, MR-beacon will be generated on that place. The initial rotation of the MR-beacon will follow the rotation of the Euler angle of the pointer during the first MR-beacon generation.
IV-D3 Move Function
IV-D4 Delete Function
IV-D5 Select Function

### v2

@@ Abstract
Recent advancements in robotics have led to the development of numerous interfaces to enhance the intuitiveness of robot navigation. However, the reliance on traditional 2D displays imposes limitations on the simultaneous visualization of information. Mixed Reality (MR) technology addresses this issue by enhancing the dimensionality of information visualization, allowing users to perceive multiple pieces of information concurrently. This paper proposes the Mixed Reality-based Robot Navigation Interface using an Optical-see-through MR-beacons (MRNaB), a novel approach that uses MR-beacons created with an “air tap”, situated in the real world. This beacon is persistent, enabling multi-destination visualization and functioning as a signal transmitter for robot navigation, eliminating the need for repeated navigation inputs. Our system is mainly constructed into four primary functions: “Add”, “Move”, “Delete”, and “Select”. These allow for the addition of MR-beacons, location movement, its deletion, and the selection of MR-beacons for navigation purposes, respectively. To validate the effectiveness, we conducted comprehensive experiments comparing MRNaB with traditional 2D navigation systems. The results show significant improvements in user performance, both objectively and subjectively, confirming that the MRNaB enhances navigation efficiency and user experience. For additional material, please check: https://mertcookimg.github.io/mrnab
@@ I Introduction
Recently, many researchers have focused on MR [7, 8, 9] as a tool for enhancing human-computer interaction by offering a more intuitive and immersive interface. MR significantly improves the user experience by providing a unified view of virtual and real-world elements, thereby minimizing the need for attention shifts and enhancing operational effectiveness[10, 11, 12]. Numerous studies have introduced innovative robot navigation interfaces through the use of an MR device called Hololens 2 [6, 13, 14, 15]. However, specifying a navigation destination within a specific location, especially during delivery processes in households or companies, can be challenging. Existing interfaces encounter limitations in two main areas: Firstly, existing interfaces use markers like arrows to set the robot’s destination, but users often struggle to visualize the robot’s status at the destination, especially in space-constrained areas. This difficulty can cause the user to fail to navigate the robot to its intended location. Secondly, users need to repeatedly specify the robot’s pose for frequently visited destinations, which diminishes efficiency.
This paper introduces MRNaB, a novel approach for robotic navigation by employing a persistent MR-beacon, which can be created by “air tap”, that mirrors the actual shape of the robot and retains its position even after the project is restarted, useful for visualization in multiple repetitive destinations. Fig. 1 shows the concept of the system. This approach not only simplifies the task of visualizing the robot’s destination in navigation but also streamlines the process of setting destinations that are regularly used. In order to implement the system, we provide the user with 4 main functions which are “Add” to add the MR-beacon from the floor, “Move” to move the location of the MR-beacon, “Select” to select the MR-beacon for robot navigation, and “Delete” to delete the MR-beacon that has been made. We also implemented a database system, which eliminates the need for users to repeatedly specify the robot’s location. This feature is particularly beneficial for environments that require repetitive navigation to the same place, such as households, especially for delivery tasks. Our contributions are summarized as follows:
  • This paper proposes MRNaB, a mixed reality navigation interface with an MR-beacon shaped like the robot, which can be used for multiple destinations by utilizing the “air tap” gesture, improving destination visualization and enhancing navigation efficiency.
@@
This paper consists of six sections including this one. Section II explains the related works. Sections III and IV introduce proposed system design and system implementation. Section V shows the experimental results to confirm the usefulness of the proposed method. Section VI concludes this paper.
@@ II Related Works
$\bullet$ Real World Spatial Understanding: A 2D map, generated by SLAM using the robot’s Lidar, provides only a rough estimation of the environment based on the Lidar’s height. This limitation becomes apparent in typical household structures, which do not conform to uniform shapes like cubes, thus complicating spatial comprehension from a 2D perspective. 3D maps, though can give the understanding of the space, but is not as clear as the real-world itself and the creation of 3D maps is resource-intensive, especially for dynamic environments with frequent changes.
Our research aims to bridge the identified challenges by augmenting user visual capabilities through the integration of real-world information and 2D maps generated by SLAM for robotic navigation. This approach addresses several key issues: it enables immediate impact on real-world robots, mitigates resource constraints, enhances spatial understanding, and eliminates the need for users to constantly shift their focus between the display and the real-world environment. This approach also adapts effectively to dynamic environments by utilizing a 2D map for robot movement and real-world information for user operations. Since the 2D map represents navigable space, small environmental changes do not require remapping, ensuring the robot can continue operating without disruption. Additionally, by enhancing spatial understanding through HoloLens 2, our system improves the user’s ability to interpret and adapt to the environment.
@@ II-B MR Robot Navigation Interface
Multi-Destination Visualization
@@
In the field of MR for robotic navigation, research initiatives have explored innovative methods to enhance human-robot interaction. Studies include using Hololens 2 to visualize robot information through occlusions [26] and employing markers and signs to indicate robot motion intent, safety distances, and interaction guidelines [27, 28]. Specialized interfaces for wheelchair robots have been developed to display intent [29, 30], and multi-user interfaces facilitate simultaneous interactions with multiple robots [31]. Lee et al. also explore the eye gaze field by using it to operate the robot manually [13]. Kot et al. combine a joystick and Hololens 2 to operate the robot manually from other places [32]. However, these studies are for manual navigation.
Walker et al.[24] proposed similar idea by using MR surrogates for teleoperating robots with a joystick, aiding users in visualizing the robot’s destination. However, their method requires additional hardware, focuses on teleoperation and only shows the last waypoint for multiple waypoints, potentially confusing users about each waypoint’s significance and its relation to the overall route, particularly in household settings.
This paper focuses on the MR interface for autonomous navigation. For autonomous navigation, one notable example is ARviz [6], which enhances 2D maps created by SLAM through the use of Hololens 2, enabling visualization of information provided by Rviz. Robot navigation is facilitated by arrows created by “air tap” hand gestures to show the goal of the navigation. Some also use waypoints for navigation goals [33, 34, 35]. Chen et al. made an approach involving an interface that employs holograms and “drag and drop” interface for visualization and navigation by 3D map created by SLAM [14]. Wu et al. use the 2D map with waypoints interface to navigate the robot [25]. Zhang et al. also explore the utilization of eye gaze and head movement for robot navigation and robot arm control [15].
@@
$\bullet$ Multi-Destination Visualization: Users are unable to visualize the actual condition or status of the robot’s destination for multiple destinations. This limitation makes navigation difficult, especially in space-constrained areas, such as household. Displaying multiple markers in a confined space can lead to confusion, as it may not be immediately clear how each marker relates to the robot’s intended movement.
@@
Our research proposes to overcome these challenges by enhancing the navigation interface by representing the robot’s goal pose with an object that matches the robot’s actual dimensions, thus making it easier for user to visualize where the robot wants to go even for multiple destinations. Our research also proposes implementing a database system that allows users to retrieve previously set destinations upon re-running the program.
To sum up, the comparison between our interface and other interfaces can be expressed in Table I. 2D interface despite having multi-destination status visualization and destination persistency, has a problem with seamless attention and real-world spatial understanding since it only uses the map. MR interface addresses this problem. However not being able to visualize the robot’s destination for multiple destinations and a feature to navigate to repetitive places make navigation difficult. This research tries to bridge all these challenges by providing all the features provided by both 2D interfaces and other MR interfaces.
Figure 2: System Design Diagram. The user interacts with Hololens 2 by using gestures to access the hand menu and interact with the MR beacon to access the beacon functions. Our system also has a database that saves the information (Pose and ID) of the robot and uses it during the first co-localization. Hololens 2 receives robot data while sending the log and goal pose to the ROS 2 side. ROS 2 has a log recorder and TF converter to convert robot and map information to only robot information. Goal pose will be sent directly using the Kachaka gprc ROS 2 bridge to the “Kachaka” robot.
@@ III System Design
In order to access the MR-beacon to navigate the robot, the hand menu is used by doing the ”Hand Constraint Palm Up” movement. Hand menu is shown in Fig. 3. In the hand menu, there will be beacon button and stage button. Stage button is only used for experiment. Beacon button will then further be divided into 6 buttons which are the back button, off button, add button, move button, select button, and delete button. Back button is used to return to the main menu and off button is used to remove all the functionality of the MR-beacon. The main functions for this system are add, move, select, and delete functions which correspond to their respective buttons.
@@
Figure 4: Add Mode Process. (a) shows the initial state where there is no MR-beacon on the floor. (b) When user does the “air tap” movement to the floor, MR-beacon will be generated and location setting starts. (c) As user drags the pointer, MR-beacon will follow the pointer. Once “air tap” is released, MR-beacon will fix the location and start the direction setting. (d) To fix the direction, user has to do one more “air tap”.
@@
Figure 5: Move Mode Process. (a) shows the initial state where there is an MR-beacon on the floor. (b) When user does the “air tap” movement to the MR-beacon, MR-beacon will be generated and the location setting starts. (c) As user drags the pointer, MR-beacon will follow the pointer. Once “air tap” is released, MR-beacon will fix the location and start the direction setting. (d) To fix the direction, user has to do one more “air tap”.
III-A Add Function
Add function is used to generate MR-beacon from the floor. After pressing add button, user will enter “Add Mode”. Fig. 4 shows the “Add Mode” process for MR-beacon.
Initially, there is no MR-beacon on the floor. When user uses “air tap” gesture (pointer down) to the floor, beacon will be generated from the floor and follow the location of the pointer of the user by dragging the pointer. In order to fixate the location of the MR-beacon, user needs to release the “air tap” and after that direction setting will start where MR-beacon direction will follow the location of the pointer of the user. To fixate the direction of the MR-beacon, user needs to do “air tap” gesture once more (pointer click).
III-B Move Function
Move function is used to move or adjust the pose of the MR-beacon to other poses. After pressing move button, user will enter “Move Mode”, and user can move the MR-beacon that was generated before. Fig. 5 shows “Move Mode” process which works similarly to “Add Mode”. However, instead of doing the “air tap” movement to the floor, the user has to do “air tap” to the MR-beacon.
The initial state starts with MR-beacon already on the floor. When user does “air tap” gesture to the MR-beacon, the MR-beacon will turn yellow and user can drag the location of the MR-beacon freely on the floor. In order to fixate the location of the MR-beacon, user needs to release the “air tap” and after that direction setting will start where MR-beacon direction will follow the location of the pointer of the user. In order to fixate the direction of the MR-beacon, user needs to do “air tap” gesture once more (pointer click).
@@ III-B Move Function
Figure 6: Select Mode Process. (a) shows the initial state where there is an MR-beacon on the floor. (b) By clicking the MR-beacon, MR-beacon’s location will be sent to the real robot. (c) Navigation will be started once robot receives the information. (d) Robot reaches the location of the MR-beacon.
@@
III-C Select Function
Select function is used to send the pose of the MR-beacon to the real robot. After pressing select button, user will enter “Select Mode”, and user can navigate the robot to any of the MR-beacon generated. “Select Mode” process is shown in Fig. 6. Initial state starts with MR-beacon on the floor already. By doing “air tap” gesture (pointer click) to the MR-beacon, the pose of the MR-beacon will be transformed to be relative to the QR code and be sent to the robot. Robot will then do navigation to the targeted pose.
III-D Delete Function
Delete function is used to delete MR-beacon. After pressing delete button, user will enter “Delete Mode” where user can delete any generated MR-beacon on the floor. “Delete Mode” process is shown in Fig. 7. Initial state starts with MR-beacon on the floor. By doing “air tap” gesture (pointer click) to any MR-beacon, it will be deleted from the floor. The information of that MR-beacon will also be deleted from the database.
@@ IV-A Communication with ROS 2
Figure 8: ROS 2 Coordinate System
Figure 9: Hololens 2 Coordinate System
@@ IV-B Co-localization
ROS 2 coordinate system (Fig. 9) follows the right-handed coordinate system where the positive $x$-axis points forward, positive $y$-axis points to the left, and positive $z$-axis points upward. Meanwhile, Hololens 2 coordinate system (Fig. 9) follows the left-handed coordinate system where the positive $x$-axis points to the right, the positive $y$-axis points upward, and the positive $z$-axis points forward.
In order to align the coordinate of ROS 2 and Hololens 2 , co-localization is required. In this system, we use the same co-localization process used by ARviz [6] by using vuforia. By using vuforia, real object’s prototype will first be put to the vuforia database and then the object will be accessed as a unity object. All of the object made in the unity system will have the parent set to this QR code. We put this QR code to be on the same location of the map location in ROS system, in order to make the transformation from Unity coordinate system to ROS coordinate system easier.
IV-C MR-beacon
IV-C1 MR-beacon Design
The dimensions of the MR-beacon are designed to mirror those of the real robot we used, as illustrated in Fig. 10. While the actual shape of the robot is not precisely a block, we simplified its form for the design, disregarding specific elements like the wheels or the camera.
IV-C2 Add Function
For MR-beacon add function, since there is no MR-beacon in the first place, we put a transparent plane object at the top of the floor whose origin is located exactly at the QR Code location. To this object, we then put object manipulator provided by MRTK, while restricting the movement to only $x$-axis and $z$-axis, as well as restricting the rotation in all directions. Once user starts doing the “air tap” movement, MR-beacon will be generated on that place. The initial rotation of the MR-beacon will follow the rotation of the Euler angle of the pointer during the first MR-beacon generation.
IV-C3 Move Function
IV-C4 Delete Function
IV-C5 Select Function
Figure 11: Database Usage in Multi-Destination Visualization and Persistency
IV-D Database
@@
The database is used to store the information about the MR-beacon. All the information of the MR-beacon that will be stored includes: First, the name of the robot which also can be used as the ID, which is created by using Globally Unique Identifier (GUID). Second, position of the MR-beacon represented by $x$, $y$, $z$ for $x$-axis, $y$-axis, and $z$-axis of the location of the robot relative to the QR code respectively. Third, rotation of the MR-beacon represented by $q_{x}$, $q_{y}$, $q_{z}$, and $q_{w}$ quaternion of the robot relative to the QR code respectively.
After the direction setting in the add function, the database will store the information of the established MR-beacon as new data. After the direction setting in the move function, the database will check the moved MR-beacon ID to update the information of the pose. After the delete function is used, the database will remove the item from the database. Lastly, when the user starts the project, if the QR code has been found, all of the stored beacons will be instantiated. The database’s pseudocode is shown in Algorithm 1. Fig. 11 illustrates the real-world implementation of the database. Each beacon retains its ID, position, and rotation, ensuring multi-destination visualization while maintaining data persistence.
V Experiment
To evaluate the effectiveness of our system, we conducted an experiment comparing our proposed MRNaB system with a traditional 2D system using a computer display and mouse, which is commonly used in daily life and, to our knowledge, represents a novel approach. The experiment was conducted twice, once for the 2D system and once for our MRNaB system. Fourteen participants (7 women and 7 men, aged 20 to 28) were involved, most of whom had no prior experience with VR/AR/MR or ROS. The order of the experiments for the conventional method and the proposed method was randomized.